OpenStudy (anonymous):

find the range of the function f(x)=x^2+8

OpenStudy (usukidoll):

the range is usually located in the y axis... since the function isn't a fraction, there is no restriction.

OpenStudy (usukidoll):

but we need to find the minimum for the range The range is different because x^2 will make any value of x either 0 or a positive number. Therefore, we need to establish the minimum let x = 0 then solve for y f(x) can also mean y y = x^2+8 if x = 0 then y=(0)^2+8